Chemical Flowsheeting and Nectar

Robert Clay Michael Hemy

Jamshid Mahdavi Todd Mummert

Carnegie Mellon University

Introduction

• Chemical Flowsheeting

• LAP solver over large cost matrices • Heterogenous Computing

Sample Generation

• Generate random samples.

• iWarp produces the samples and sends via HIPPI to the C90

• HIB physically connected at 4 points in the array

• Results 1. iWarp: 1450K samples/sec

2. alpha: 185K

3. Sun 4/330: 20K samples/sec

Sending Data over Hippi

• iWarp → C90 — Michael Hemy’s HIB routines.

• C90 ↔ CM2 — PSC supplied DHSC routines.

1. Advantages: synchronization, data conversion, moderation

2. Disadvantages: inflexible, high latency due to PVM

Application Results

• Successful completion on small cost matrices.

• Computation limited on iWarp (6MB/s)

• Problems w/ larger cost matrices 1. Insufficient space on C90.

2. Pipelining

Current Plans

• Remove DHSC dependencies • Change Data Flow